an equation that goes through the point (1, 5) and has a slope of 2/3
Answer by jim_thompson5910(35256)

y-y1 = m(x-x1)


y-5=(2/3)(x-1)


y-5=(2/3)x-(2/3)(1)


y-5=(2/3)x-2/3


y=(2/3)x-2/3+5


y=(2/3)x+13/3


So the answer is


If you need more help, email me at jim_thompson5910@hotmail.com

Also, please consider visiting my website: http://www.freewebs.com/jimthompson5910/home.html and making a donation. Thank you

Jim